1. Prednisone 10 mg/kg/day is ordered for a child weighing 25lbs. Prednisone is available in 5 mg/5ml.

a. How many mg per dose will the child receive?

b. How many mg per day will the child receive?

c. How many ml per dose will the child receive?

2 Order: Ferrous sulfate 9 mg po tid. child weighs 13lbs. The safe dosage is 3-6 mg/kg/24hrs

a. what is the safe 24 hours range?

b is the ordered dose safe for a 24 hr period

3. Order Cefaclor 180 mg po q 8hr. child weighs 10lbs. The safe dose is 15 mg/kg dose.

a what is the safe dose for this child

b is the ordered dose safe

c if the medication is supplied 125mg/5ml, how many ml will you administer
Answer #1
  1. ordered Prednisone= 10 mg/kg/day for a child weighing 25lbs= 11.36 kg

Available Prednisone = 5 mg/5ml.

a. The amount in mg per dose the child will receive=10 X 11.36=113.6 mg as there is no description od dose like OD, BID or t.i.d

b. The amount in mg per day will the child receive will be= 113.6 as this is per day dose

c. The dose in ml will be =?

Given that 5mg/5ml

Means 1mg/ml

Hence for 113.6 mg it will be 113.6 ml

  1. Ordered Ferrous sulfate= 9 mg po tid.

Child weight= 13lbs=5.90 kg

The safe dosage is 3-6 mg/kg/24hrs

a. The safe dose range for 24 hours in this case will be=17.7- 35.4mg

b. The ordered dose i.e. 9 mg po tid or 27mg for 24 hours is safe dose as it is within the safe dose range i.e. 17.7- 35.4mg for a 24 hr period

  1. Ordered Cefaclor = 180 mg po q 8hr

Child weight= 10lbs =4.54 kg

The safe dose is 15 mg/kg/dose.

a The safe dose for this child will be =4.54X 15= 68.1 mg/dose

b The ordered dose is not safe as it is above the safe dose level

c If the medication is supplied 125mg/5ml

Means 25mg/ml

For 180 mg required ml= 180/25=7.2 ml the nurse should administer per dose as per the order